5)Garlington Technologies Inc.'s 2019 financial statements are shown below:

Income Statement for December 31, 2019

Sales $4,000,000
Operating costs 3,200,000
EBIT $  800,000
Interest 120,000
Pre-tax earnings $  680,000
Taxes (25%) 170,000
Net income 510,000
Dividends $  190,000

Balance Sheet as of December 31, 2019

Cash $   160,000 Accounts payable $   360,000
Receivables 360,000 Line of credit 0
Inventories 720,000 Accruals 200,000
Total CA $1,240,000 Total CL $   560,000
Fixed assets 4,000,000 Long-term bonds 1,000,000
Total Assets $5,240,000 Common stock 1,100,000
RE 2,580,000
Total L&E $5,240,000

Suppose that in 2020 sales increase to $4.6 million and that 2020 dividends will increase to $200,000. Forecast the financial statements using the forecasted financial statement method. Assume the firm operated at full capacity in 2019. The long-term bonds have an interest rate of 11%. New financing will be with a line of credit. Assume it will be added at the end of the year. Cash does not earn any interest income. Enter your answers as positive values. Do not round intermediate calculations. Round your answers to the nearest dollar.

Garlington Technologies Inc.
Pro Forma Income Statement
December 31, 2020
Sales $  
Operating costs $  
EBIT $  
Interest $  
Pre-tax earnings $  
Taxes (25%) $  
Net income $  
Dividends: $  
Addition to RE: $  
Garlington Technologies Inc.
Pro Forma Balance Statement
December 31, 2020
Cash $  
Receivables $  
Inventories $  
Total current assets $  
Fixed assets $  
Total assets $  
Accounts payable $  
Line of credit $  
Accruals $  
Total current liabilities $  
LT bonds $  
Common stock $  
Retained earnings $  
Total L&E $  
Answer #1

Growth rate in sales = g = 4.6 / 4.0 - 1 = 15% (the sales figures are in $ million)

Please see the table below. Please see the linkage column to understand the mathematics. Line of credit is the balancing figure and will be calculated last after all other items have been calculated.

Garlington Technologies Inc.
Pro Forma Income Statement
31-Dec-20 Linkage $
Sales A = 4.6 million    4,600,000
Operating costs B = 80% x A    3,680,000
EBIT C = A - B       920,000
Interest D = 11% x 1,000,000 of debt       110,000
Pre-tax earnings E = C - D       810,000
Taxes (25%) F = 25% x E       202,500
Net income G = E - F       607,500
Dividends: H       200,000
Addition to RE: I = G - H       407,500
Pro Forma Balance Statement
31-Dec-20
Cash J = 160,000 x (1 + 15%)       184,000
Receivables K = 360,000 x (1 + 15%)       414,000
Inventories L = 720,000 x (1 + 15%)       828,000
Total current assets M = J + K + L 1,426,000
Fixed assets N = 4,000,000 x (1 + 15%)    4,600,000
Total assets O = M + N 6,026,000
Accounts payable P = 360,000 x (1 + 15%)       414,000
Line of credit Q = Balancing figure       294,500
Accruals R = 200,000 x (1 + 15%)       230,000
Total current liabilities S = P + Q + R       938,500
LT bonds T = last year figure    1,000,000
Common stock U = Last year figure    1,100,000
Retained earnings V = 2,580,000 + I calculated above    2,987,500
Total L&E W = S + T + U + V 6,026,000